OpenStudy (anonymous):

Find the product: (8x^2 – 5y) (8x^2 + 5y)

OpenStudy (anonymous):

FOIL

OpenStudy (anonymous):

i know to use foil but i dont understand how

OpenStudy (anonymous):

constants multiply together, as for like variables with exponents >>> x^y * x^z = x^y+z

OpenStudy (anonymous):

so the first part in the FOIL operation is (8x^2)(8x^2) = 64x^4

OpenStudy (anonymous):

ok so when it comes time to put them together do i subtract? for instance the 64x^4 - (number) ?

OpenStudy (anonymous):

depends, the second part is (8x^2)(5y) which is a positive number so you would add that. the third part however, (-5y)(8x^2) is a negative number so you subtract

OpenStudy (anonymous):

thank you! that confused me but i think that makes a little more sense!

OpenStudy (anonymous):

np